# Numerical stability of force-gradient integrators and their Hessian-free variants in lattice QCD simulations

### v1

@@ 3 Linear stability of force-gradient integrators
The force-gradient term (10) for the harmonic oscillator (15b) reads $\hat{\mathcal{C}}=2\omega^{3}q\frac{\partial}{\partial p}$. A decomposition algorithm will typically become unstable for $\lvert z\rvert>z_{*}$, where $z\coloneqq\omega{h}$ and the parameter $z_{*}$ represents the stability threshold of the integrator. The objective is to determine the parameter $z_{*}$ for a given force-gradient integrator by adapting the theory for splitting methods as presented in [BlanesStability]. Starting from the initial values $(q(0),p(0))=(q_{0},p_{0})$, the exact solution to the harmonic oscillator (15b) at time point ${h}>0$ is given by
@@
$\begin{split}\mathbf{K}(z)&=\prod_{\ell=1}^{s}\begin{pmatrix}1&0\\ -b_{s+1-\ell}z+2c_{s+1-\ell}z^{3}&1\end{pmatrix}\begin{pmatrix}1&a_{s+1-\ell}z\\ 0&1\end{pmatrix},\end{split}$ (16)

### v2

@@ 3 Linear stability of force-gradient integrators
The force-gradient term (10) for the harmonic oscillator (15b) reads $\hat{\mathcal{C}}=2\omega^{3}q\frac{\partial}{\partial p}$. A decomposition algorithm will typically become unstable for $\lvert z\rvert>z_{*}$, where $z\coloneqq\omega{h}$ and the parameter $z_{*}$ represents the stability threshold of the integrator. The objective is to determine the parameter $z_{*}$ for a given force-gradient integrator by adapting the theory for splitting methods as presented in BlanesStability . Starting from the initial values $(q(0),p(0))=(q_{0},p_{0})$, the exact solution to the harmonic oscillator (15b) at time point h>0{h}>0 is given by
@@
[…]